# Stability Analysis of Grünwald Interpolation Operators on Chebyshev Nodes

### v1

@@ Abstract
In 1941, G. Grünwald proved the convergence of a sequence of operators constructed using classical Lagrange interpolation at Chebyshev nodes. In this work, we establish a perturbed version of Grünwald’s result, thereby extending the class of admissible nodal points. Specifically, we provide sufficient conditions for convergence when the interpolation nodes are of the form $\{\cos\eta_{k}\}_{k=1}^{n}$, where $\{\eta_{k}\}_{k=1}^{n}$ is a general sequence. We refer to these operators as Grünwald operators. In particular, we prove a convergence result when $\{\eta_{k}\}_{k=1}^{n}$ is equidistant and uniformly distributed. We establish a Voronovskaja-type estimate for the convergence of these operators and derive quantitative results involving the first and second moduli of smoothness.
@@ 1 Introduction
The aim of this article is to investigate for which sets of nodal points the sequence of operators $\{\mathcal{G}_{n}\}$ converges. As Grünwald has already resolved this problem in the case of Chebyshev nodes, our objective is to extend this result to other choices of nodal points. This article is structured as follows. In the following section, we address this problem by proving a perturbed version of Theorem 1.3, thereby identifying a broader class of nodal points for which the convergence result holds. Specifically, we establish the result for the case where the $\eta_{k}$’s in the Definition 1.2 are uniformly distributed and equidistant. We also present an intermediate result that emerges naturally from the proof of our main lemma. Following this, we prove a Voronovskaja-type estimate for the convergence of $\{\mathcal{G}_{n}\}$. In the subsequent section, we obtain quantitative estimates for the convergence of $\{\mathcal{G}_{n}\}$ using first and second modulus of smoothness.
@@ 10 $0,\frac{\pi}{2}],(\frac{\pi}{2},\pi]$ $\frac{\pi}{2} >$0>0 $\leq 2$
$\eta_{0}=(\eta_{0}-\frac{\pi}{2n})+\theta_{1}^{(n)}$
$\eta_{1}+\beta_{n}=\eta_{0}+\frac{\pi}{n}=(\eta_{0}-\frac{\pi}{2n})+\theta_{2}^{(n)}$
$\eta_{2}+2\beta_{n}=\eta_{0}+\frac{2\pi}{n}=(\eta_{0}-\frac{\pi}{2n})+\theta_{3}^{(n)}$
@@
We have, $\eta_{0}\leq\frac{\pi}{n}$, so that $\eta_{0}-\frac{\pi}{2n}\in(-\frac{\pi}{2n},\frac{\pi}{2n})$. Now, it is easy to apply the Lemma 2.4 by choosing $\theta_{0}=\eta_{0}-\frac{\pi}{2n}$. Hence the result follows. ∎
@@
In the following section, we obtain a Voronovskaja-type estimate for the convergence of the sequence $\{\mathcal{G}_{n}\}$. Thereafter, we derive quantitative estimates for the rate of convergence of $\{\mathcal{G}_{n}\}$ using the first and second moduli of smoothness. Throughout this analysis, the nodes are taken to be the perturbed nodes $\{\cos\tilde{\theta}_{k}^{(n)}\}_{k=1}^{n}$ as described in Lemma 2.4.
@@ 3 A Voronovskaja-type estimate
We now prove a Voronovskaja-type estimate for the sequence $\{\mathcal{G}_{n}\}$. A full Voronovskaja-type theorem cannot be obtained in this setting; nevertheless, we derive an estimate for the convergence. For this purpose, we obtain the following lemma.
Lemma 3.1.
@@
$\begin{split}\frac{\sin\theta\,\sin\tilde{\theta}_{1}^{(n)}}{\sin\frac{1}{2}\left(\theta+\tilde{\theta}_{1}^{(n)}+\frac{\pi}{2n}\right)\cdot\sin\frac{1}{2}\left(\theta+\tilde{\theta}_{1}^{(n)}-\frac{\pi}{2n}\right)}&=\frac{\sin\theta\,\sin\left(\frac{\pi}{2n}-\theta_{0}\right)}{\sin\frac{1}{2}\left(\theta+\frac{\pi}{n}-\theta_{0}\right)\cdot\sin\frac{1}{2}(\theta-\theta_{0})}\\[6.0pt] &=\frac{2\cos\frac{\theta}{2}\sin\frac{\theta}{2}\,\sin\left(\frac{\pi}{2n}-\theta_{0}\right)}{\sin\frac{1}{2}\left(\theta+\frac{\pi}{n}-\theta_{0}\right)\sin\frac{1}{2}(\theta-\theta_{0})}.\end{split}$
For $\theta\leq\frac{\pi}{2}$, we have $\sin\frac{\theta}{2}\leq\sin\frac{1}{2}\left(\theta+\frac{\pi}{n}-\theta_{0}\right)\quad\text{and}\quad\sin\left(\frac{\pi}{2n}-\theta_{0}\right)\leq\sin\frac{1}{2}(\theta-\theta_{0}),$ since $\kappa_{n}>\frac{\pi}{2n}$.
For $\theta>\frac{\pi}{2}$, we have $\cos\frac{\theta}{2}\leq\sin\frac{1}{2}\left(\theta+\frac{\pi}{n}-\theta_{0}\right)=\cos\frac{1}{2}\left(\pi-\theta-\frac{\pi}{n}+\theta_{0}\right),$ since $\cos\theta$ is non-increasing.
Note that $\frac{\theta}{2}\geq\frac{1}{2}\left(\pi-\theta-\frac{\pi}{n}+\theta_{0}\right).$ holds since $\theta>\frac{\pi}{2}\geq\frac{\pi}{2}-\frac{\pi}{2n}+\frac{\theta_{0}}{2}.$ In either case, inequality 3.2 holds.
Case 2: k=n $|\theta-\tilde{\theta}_{n}^{(n)}|>\kappa_{n}$. In this case, either $\theta>\frac{(2n-1)\pi}{2n}-\theta_{0}+\kappa_{n}\quad\text{or}\quad\theta<\frac{(2n-1)\pi}{2n}-\theta_{0}-\kappa_{n}.$ Since $\kappa_{n}>\frac{\pi}{2n}$, the first case cannot occur. Hence, $\theta<\frac{(2n-1)\pi}{2n}-\theta_{0}-\kappa_{n}=\pi-\frac{\pi}{n}-\theta_{0}-\kappa_{n}.$
@@
For all $\theta\in[0,\pi]$, since $\cos\theta$ is non increasing, $\cos\frac{\theta}{2}\leq\cos\frac{1}{2}(\theta-\theta_{0}).$
Let $\theta\leq\frac{\pi}{2}$. Since $\cos\theta$ is increasing on $0,\frac{\pi}{2}]$, we have $\sin\frac{\theta}{2}=\cos\Big(\frac{\pi}{2}-\frac{\theta}{2}\Big)\leq\cos\frac{1}{2}\big(\theta-\frac{\pi}{n}-\theta_{0}\big),$ since $\frac{\pi}{2}-\frac{\theta}{2}>\frac{1}{2}(\theta-\frac{\pi}{n}-\theta_{0}).$
For $\theta>\frac{\pi}{2}$, we have $\sin\frac{\theta}{2}=\cos\Big(\frac{\pi}{2}-\frac{\theta}{2}\Big).$ Since $\cos\theta$ is non increasing on $0,\pi]$, $\cos\Big(\frac{\pi}{2}-\frac{\theta}{2}\Big)\leq\cos\frac{1}{2}(\theta-\frac{\pi}{n}-\theta_{0}).$ Now, $\frac{\pi}{2}-\frac{\theta}{2}\geq\frac{1}{2}(\theta-\frac{\pi}{n}-\theta_{0}),$ which holds if and only if $\theta<3\theta_{0}+\pi-\frac{\pi}{n},$ which is true since $\theta<\pi-\frac{\pi}{n}-\theta_{0}-\kappa_{n}.$ Hence by the above cases, 3.2 is verified. ∎
@@
$\limsup\limits_{n\to\infty}n^{\frac{1}{3}}|\mathcal{G}_{n}(f)(\theta)-f(\theta)|\leq(\frac{9\pi^{4}}{64}+c_{2})|f^{\prime}(\theta)|+\frac{9\pi^{5}}{128}|f^{\prime\prime}(\theta)|+\frac{9M\pi^{5}}{64},$
where $M=\|f^{\prime\prime}\|_{\infty}$ and $c_{2}$ is as in Lemma 2.4.
@@
Let $f\in C^{2}[0,\pi]$, $t,\theta\in[0,\pi]$. Then by the Taylor’s formula we have
$f(t)-f(\theta)=f^{\prime}(\theta)(t-\theta)+f^{\prime\prime}(\theta)\frac{(t-\theta)^{2}}{2}+(f^{\prime\prime}(c)-f^{\prime\prime}(\theta))\frac{(t-\theta)^{2}}{2},$
where $c$ lies between $t$ and $\theta$. By Lemma 2.4, we have,
$\frac{1}{2}\sum\limits_{k=1}^{n}|l_{k}(\theta-\frac{\pi}{2n})+l_{k}(\theta+\frac{\pi}{2n})|\leq c_{2}.$ (3.3)
Let $e_{0}(t)=t$ for all $t\in[0,\pi]$ and let $\epsilon_{n}=\frac{1}{n^{\frac{1}{3}}}$ for $n=1,2,\ldots$ . Then
$\begin{split}|\mathcal{G}_{n}(e_{0})(\theta)-e_{0}(\theta)|&\leq\frac{1}{2}\sum\limits_{1\leq k\leq n,|\theta-\tilde{\theta}_{k}^{(n)}|<\frac{\pi}{2n}+\epsilon_{n}}|\tilde{\theta}_{k}^{(n)}-\theta||l_{k}(\theta-\frac{\pi}{2n})+l_{k}(\theta+\frac{\pi}{2n})|\\ &\quad\quad\quad+\frac{1}{2}\sum\limits_{1\leq k\leq n,|\theta-\tilde{\theta}_{k}^{(n)}|\geq\frac{\pi}{2n}+\epsilon_{n}}|\tilde{\theta}_{k}^{(n)}-\theta||l_{k}(\theta-\frac{\pi}{2n})+l_{k}(\theta+\frac{\pi}{2n})|\\ &\leq(\frac{\pi}{2n}+\epsilon_{n})c_{2}+\pi\times\frac{9\pi^{3}}{64n^{2}}\sum\limits_{k=1}^{n}\frac{1}{\left(\theta-\tilde{\theta}_{k}^{(n)}-\frac{\pi}{2n}\right)^{2}}\\ &\leq(\frac{\pi}{2n}+\epsilon_{n})c_{2}+\pi\times\frac{9\pi^{3}}{64n^{2}}\sum\limits_{k=1}^{n}\frac{1}{\epsilon_{n}^{2}}\\ &=(\frac{\pi}{2n}+\epsilon_{n})c_{2}+\frac{9\pi^{4}}{64}\frac{1}{n\epsilon_{n}^{2}}.\end{split}$
In the first sum we use 3.3 and in the second sum use the fact that $|\theta_{k}^{(n)}-\theta|\leq\pi$ and Lemma 3.1 where $\kappa_{n}=\frac{\pi}{2n}+\epsilon_{n}>\frac{\pi}{2n}$. Similarly, we have
@@
$\begin{split}|\mathcal{G}_{n}((e_{0}-\theta)^{2})|&\leq\frac{1}{2}\sum\limits_{1\leq k\leq n,|\theta-\tilde{\theta}_{k}^{(n)}|<\frac{\pi}{2n}+\epsilon_{n}}(\tilde{\theta}_{k}^{(n)}-\theta)^{2}|l_{k}(\theta-\frac{\pi}{2n})+l_{k}(\theta+\frac{\pi}{2n})|\\ &\quad\quad\quad+\frac{1}{2}\sum\limits_{1\leq k\leq n,|\theta-\tilde{\theta}_{k}^{(n)}|\geq\frac{\pi}{2n}+\epsilon_{n}}(\tilde{\theta}_{k}^{(n)}-\theta)^{2}|l_{k}(\theta-\frac{\pi}{2n})+l_{k}(\theta+\frac{\pi}{2n})|\\ &\leq(\frac{\pi}{2n}+\epsilon_{n})^{2}c_{2}+\pi^{2}\times\frac{9\pi^{3}}{64n^{2}}\sum\limits_{k=1}^{n}\frac{1}{\epsilon_{n}^{2}}\\ &\leq(\frac{\pi}{2n}+\epsilon_{n})^{2}c_{2}+\frac{9\pi^{5}}{64}\frac{1}{n\epsilon_{n}^{2}}\\ &=(\frac{\pi^{2}}{4n^{2}}+\epsilon_{n}^{2}+\frac{\pi\epsilon_{n}}{n})c_{2}+\frac{9\pi^{5}}{64}\frac{1}{n\epsilon_{n}^{2}}.\end{split}$
Suppose that $\|f^{\prime\prime}\|_{\infty}=M$. Then we also have,
$\begin{split}|\mathcal{G}_{n}((e_{0}-\theta)^{2}|f^{\prime\prime}(c)-f^{\prime\prime}(\theta)|)|&\leq|f^{\prime\prime}(c)-f^{\prime\prime}(\theta)||\mathcal{G}_{n}((e_{0}-\theta)^{2})|\\ &\leq 2M\{(\frac{\pi^{2}}{4n^{2}}+\epsilon_{n}^{2}+\frac{\pi\epsilon_{n}}{n})c_{2}+\frac{9\pi^{5}}{64}\frac{1}{n\epsilon_{n}^{2}}\}.\end{split}$
Now,
[…]
$n\epsilon_{n}^{2}|\mathcal{G}_{n}((e_{0}-\theta)^{2}|f^{\prime\prime}(c)-f^{\prime\prime}(\theta)|)|\leq 2M\{(\frac{\pi^{2}\epsilon_{n}^{2}}{4n}+n\epsilon_{n}^{4}+{\pi\epsilon_{n}^{3}})c_{2}+\frac{9\pi^{5}}{64}\}.$
Since $\epsilon_{n}=\frac{1}{n^{\frac{1}{3}}}$ for $n=1,2,\ldots$, we have $\epsilon_{n},n\epsilon_{n}^{4},\frac{1}{n\epsilon_{n}^{2}}\to 0$ as $n\to\infty$ and $n\epsilon_{n}^{3}=1$.
Thus we have
@@
$\limsup\limits_{n\to\infty}n^{\frac{1}{3}}|\mathcal{G}_{n}(f)(\theta)-f(\theta)|\leq(\frac{9\pi^{4}}{64}+c_{2})|f^{\prime}(\theta)|+\frac{9\pi^{5}}{128}|f^{\prime\prime}(\theta)|+\frac{9M\pi^{5}}{64}.$

### v2

@@ Abstract
In 1941, G. Grünwald proved the convergence of a sequence of operators constructed using classical Lagrange interpolation at Chebyshev nodes. In this work, we establish a perturbed version of Grünwald’s result, thereby extending the class of admissible nodal points. Specifically, we provide sufficient conditions for convergence when the interpolation nodes are of the form $\{\cos\eta_{k}\}_{k=1}^{n}$, where $\{\eta_{k}\}_{k=1}^{n}$ is a general sequence. We refer to these operators as Grünwald operators. In particular, we prove a convergence result when $\{\eta_{k}\}_{k=1}^{n}$ is equidistant and uniformly distributed. We establish a Voronovskaja-type estimate for the convergence of these operators and derive quantitative results using modulus of continuity.
@@ 1 Introduction
The aim of this article is to investigate for which sets of nodal points the sequence of operators $\{\mathcal{G}_{n}\}$ converges. As Grünwald has already resolved this problem in the case of Chebyshev nodes, our objective is to extend this result to other choices of nodal points. This article is structured as follows. In the following section, we address this problem by proving a perturbed version of Theorem 1.3, thereby identifying a broader class of nodal points for which the convergence result holds. Specifically, we establish the result for the case where the $\eta_{k}$’s in the Definition 1.2 are uniformly distributed and equidistant. We also present an intermediate result that emerges naturally from the proof of our main lemma. Following this, we prove a Voronovskaja-type estimate for the convergence of $\{\mathcal{G}_{n}\}$. In the subsequent section, we obtain quantitative estimates for the convergence of $\{\mathcal{G}_{n}\}$ using modulus of continuity.
@@ 10 $0,\frac{\pi}{2}],(\frac{\pi}{2},\pi]$ $\frac{\pi}{2} >$0>0 $\leq 2$
$\eta_{0}=(\eta_{0}-\frac{\pi}{2n})+\theta_{1}^{(n)},$
$\eta_{1}+\beta_{n}=\eta_{0}+\frac{\pi}{n}=(\eta_{0}-\frac{\pi}{2n})+\theta_{2}^{(n)},$
$\eta_{2}+2\beta_{n}=\eta_{0}+\frac{2\pi}{n}=(\eta_{0}-\frac{\pi}{2n})+\theta_{3}^{(n)},$
@@
We have, $\eta_{0}\leq\frac{\pi}{n}$, so that $\eta_{0}-\frac{\pi}{2n}\in(-\frac{\pi}{2n},\frac{\pi}{2n})$. Now, it is easy to apply Lemma 2.4 by choosing $\theta_{0}=\eta_{0}-\frac{\pi}{2n}$. Hence the result follows. ∎
@@
In the following section, we obtain a Voronovskaja-type estimate for the convergence of the sequence $\{\mathcal{G}_{n}\}$. Thereafter, we derive quantitative estimates for the rate of convergence of $\{\mathcal{G}_{n}\}$ using the first modulus of smoothness. Throughout this analysis, the nodes are taken to be the perturbed nodes $\{\cos\tilde{\theta}_{k}^{(n)}\}_{k=1}^{n}$ as described in Lemma 2.4.
@@ 3 A Voronovskaja-type estimate
We now prove a Voronovskaja-type estimate for the sequence $\{\mathcal{G}_{n}\}$. A full Voronovskaja-type theorem cannot be obtained in this setting; nevertheless, we derive an estimate for the convergence. For this purpose, we obtain the following.
Proposition 3.1.
@@
$\begin{split}\frac{\sin\theta\,\sin\tilde{\theta}_{1}^{(n)}}{\sin\frac{1}{2}\left(\theta+\tilde{\theta}_{1}^{(n)}+\frac{\pi}{2n}\right)\cdot\sin\frac{1}{2}\left(\theta+\tilde{\theta}_{1}^{(n)}-\frac{\pi}{2n}\right)}&=\frac{\sin\theta\,\sin\left(\frac{\pi}{2n}-\theta_{0}\right)}{\sin\frac{1}{2}\left(\theta+\frac{\pi}{n}-\theta_{0}\right)\cdot\sin\frac{1}{2}(\theta-\theta_{0})}\\[5.69054pt] &=\frac{2\cos\frac{\theta}{2}\sin\frac{\theta}{2}\,\sin\left(\frac{\pi}{2n}-\theta_{0}\right)}{\sin\frac{1}{2}\left(\theta+\frac{\pi}{n}-\theta_{0}\right)\sin\frac{1}{2}(\theta-\theta_{0})}.\end{split}$
For $\theta\leq\frac{\pi}{2}$, we have
$\sin\frac{\theta}{2}\leq\sin\frac{1}{2}\left(\theta+\frac{\pi}{n}-\theta_{0}\right),\quad\text{and}\quad\sin\left(\frac{\pi}{2n}-\theta_{0}\right)\leq\sin\frac{1}{2}(\theta-\theta_{0}),$
since $\kappa_{n}>\frac{\pi}{2n}$.
For $\theta>\frac{\pi}{2}$, we have
$\cos\frac{\theta}{2}\leq\sin\frac{1}{2}\left(\theta+\frac{\pi}{n}-\theta_{0}\right)=\cos\frac{1}{2}\left(\pi-\theta-\frac{\pi}{n}+\theta_{0}\right),$
since $\cos\theta$ is non-increasing.
Note that
$\frac{\theta}{2}\geq\frac{1}{2}\left(\pi-\theta-\frac{\pi}{n}+\theta_{0}\right),$
which holds since
$\theta>\frac{\pi}{2}\geq\frac{\pi}{2}-\frac{\pi}{2n}+\frac{\theta_{0}}{2}.$
In either case, inequality 3.2 holds.
Case 2: k=n $|\theta-\tilde{\theta}_{n}^{(n)}|>\kappa_{n}$.
In this case, either
$\theta>\frac{(2n-1)\pi}{2n}-\theta_{0}+\kappa_{n}\quad\text{or}\quad\theta<\frac{(2n-1)\pi}{2n}-\theta_{0}-\kappa_{n}.$
Since $\kappa_{n}>\frac{\pi}{2n}$, the first case cannot occur. Hence,
$\theta<\frac{(2n-1)\pi}{2n}-\theta_{0}-\kappa_{n}=\pi-\frac{\pi}{n}-\theta_{0}-\kappa_{n}.$
@@
For all $\theta\in[0,\pi]$, since $\cos\theta$ is non-increasing,
$\cos\frac{\theta}{2}\leq\cos\frac{1}{2}(\theta-\theta_{0}).$
Let $\theta\leq\frac{\pi}{2}$. Since $\cos\theta$ is increasing on $0,\frac{\pi}{2}]$, we have
$\sin\frac{\theta}{2}=\cos\Big(\frac{\pi}{2}-\frac{\theta}{2}\Big)\leq\cos\frac{1}{2}\big(\theta-\frac{\pi}{n}-\theta_{0}\big),$
since
$\frac{\pi}{2}-\frac{\theta}{2}>\frac{1}{2}(\theta-\frac{\pi}{n}-\theta_{0}).$
For $\theta>\frac{\pi}{2}$, we have
$\sin\frac{\theta}{2}=\cos\Big(\frac{\pi}{2}-\frac{\theta}{2}\Big).$
Since $\cos\theta$ is non-increasing on $0,\pi]$,
$\cos\Big(\frac{\pi}{2}-\frac{\theta}{2}\Big)\leq\cos\frac{1}{2}(\theta-\frac{\pi}{n}-\theta_{0}).$
Now,
$\frac{\pi}{2}-\frac{\theta}{2}\geq\frac{1}{2}(\theta-\frac{\pi}{n}-\theta_{0}),$
which holds if and only if $\theta<3\theta_{0}+\pi-\frac{\pi}{n},$ which is true since $\theta<\pi-\frac{\pi}{n}-\theta_{0}-\kappa_{n}.$
Hence, by the above cases, 3.2 is verified. ∎
Let $c_{2}$ be the constant in the equation 2.
@@
$\limsup_{n\to\infty}n^{1/3}\,|\mathcal{G}_{n}(f)(\theta)-f(\theta)|\leq\left(\frac{9\pi^{4}}{64}+c_{2}\right)|f^{\prime}(\theta)|+\frac{9\pi^{5}}{128}|f^{\prime\prime}(\theta)|+\frac{9M\pi^{5}}{64},$
where $M=\|f^{\prime\prime}\|_{\infty}$.
@@
Let $f\in C^{2}[0,\pi]$ and $t,\theta\in[0,\pi]$. By Taylor’s theorem,
$f(t)-f(\theta)=f^{\prime}(\theta)(t-\theta)+\frac{f^{\prime\prime}(\theta)}{2}(t-\theta)^{2}+\frac{f^{\prime\prime}(c)-f^{\prime\prime}(\theta)}{2}(t-\theta)^{2},$
where $c$ lies between $t$ and $\theta$.
From equation 2, we have
$\frac{1}{2}\sum_{k=1}^{n}\bigl|l_{k}(\theta-\tfrac{\pi}{2n})+l_{k}(\theta+\tfrac{\pi}{2n})\bigr|\leq c_{2}.$ (3.3)
Let $e_{0}(t)=t$ for $t\in[0,\pi]$ and define $\epsilon_{n}=n^{-1/3}$. Then
@@
$|\mathcal{G}_{n}(e_{0})(\theta)-e_{0}(\theta)|$ $\leq\frac{1}{2}\sum_{\begin{subarray}{c}1\leq k\leq n\\ |\theta-\tilde{\theta}_{k}^{(n)}|<\frac{\pi}{2n}+\epsilon_{n}\end{subarray}}|\tilde{\theta}_{k}^{(n)}-\theta|\bigl|l_{k}(\theta-\tfrac{\pi}{2n})+l_{k}(\theta+\tfrac{\pi}{2n})\bigr|$
$\quad+\frac{1}{2}\sum_{\begin{subarray}{c}1\leq k\leq n\\ |\theta-\tilde{\theta}_{k}^{(n)}|\geq\frac{\pi}{2n}+\epsilon_{n}\end{subarray}}|\tilde{\theta}_{k}^{(n)}-\theta|\bigl|l_{k}(\theta-\tfrac{\pi}{2n})+l_{k}(\theta+\tfrac{\pi}{2n})\bigr|$
$\leq\left(\frac{\pi}{2n}+\epsilon_{n}\right)c_{2}+\pi\cdot\frac{9\pi^{3}}{64n^{2}}\sum_{k=1}^{n}\frac{1}{(\theta-\tilde{\theta}_{k}^{(n)}-\frac{\pi}{2n})^{2}}$
$\leq\left(\frac{\pi}{2n}+\epsilon_{n}\right)c_{2}+\frac{9\pi^{4}}{64}\frac{1}{n\epsilon_{n}^{2}},$
using (3.3), $|\theta-\tilde{\theta}_{k}^{(n)}|\leq\pi$, and Proposition 3.1 with $\kappa_{n}=\frac{\pi}{2n}+\epsilon_{n}>\frac{\pi}{2n}$.
Similarly,
$|\mathcal{G}_{n}((e_{0}-\theta)^{2})(\theta)|$ $\leq\left(\frac{\pi}{2n}+\epsilon_{n}\right)^{2}c_{2}+\frac{9\pi^{5}}{64}\frac{1}{n\epsilon_{n}^{2}}$
$=\left(\frac{\pi^{2}}{4n^{2}}+\epsilon_{n}^{2}+\frac{\pi\epsilon_{n}}{n}\right)c_{2}+\frac{9\pi^{5}}{64}\frac{1}{n\epsilon_{n}^{2}}.$
Let $M=\|f^{\prime\prime}\|_{\infty}$. Then
$|\mathcal{G}_{n}\big((e_{0}-\theta)^{2}(f^{\prime\prime}(c)-f^{\prime\prime}(\theta))\big)(\theta)|\leq 2M\left\{\left(\frac{\pi^{2}}{4n^{2}}+\epsilon_{n}^{2}+\frac{\pi\epsilon_{n}}{n}\right)c_{2}+\frac{9\pi^{5}}{64}\frac{1}{n\epsilon_{n}^{2}}\right\}.$
Now multiply these inequalities by $n\epsilon_{n}^{2}$:
$n\epsilon_{n}^{2}|\mathcal{G}_{n}(e_{0})(\theta)-e_{0}(\theta)|\leq\left(\frac{\pi\epsilon_{n}^{2}}{2}+n\epsilon_{n}^{3}\right)c_{2}+\frac{9\pi^{4}}{64},$
$n\epsilon_{n}^{2}|\mathcal{G}_{n}((e_{0}-\theta)^{2})(\theta)|\leq\left(\frac{\pi^{2}\epsilon_{n}^{2}}{4n}+n\epsilon_{n}^{4}+\pi\epsilon_{n}^{3}\right)c_{2}+\frac{9\pi^{5}}{64},$
$n\epsilon_{n}^{2}\big|\mathcal{G}_{n}((e_{0}-\theta)^{2}(f^{\prime\prime}(c)-f^{\prime\prime}(\theta)))(\theta)\big|\leq 2M\left(\frac{\pi^{2}\epsilon_{n}^{2}}{4n}+n\epsilon_{n}^{4}+\pi\epsilon_{n}^{3}\right)c_{2}+\frac{18M\pi^{5}}{64}.$
Since $\epsilon_{n}=n^{-1/3}$, we have
$\epsilon_{n}\to 0,\qquad n\epsilon_{n}^{4}\to 0,\qquad\frac{1}{n\epsilon_{n}^{2}}\to 0,\quad as\ n\to\infty,\qquad n\epsilon_{n}^{3}=1.$
Combining the estimates from the Taylor expansion shows that
@@
$\limsup_{n\to\infty}n^{1/3}\,|\mathcal{G}_{n}(f)(\theta)-f(\theta)|\leq\left(\frac{9\pi^{4}}{64}+c_{2}\right)|f^{\prime}(\theta)|+\frac{9\pi^{5}}{128}|f^{\prime\prime}(\theta)|+\frac{9M\pi^{5}}{64},$
as claimed. ∎